OpenStudy (anonymous):

How do you simplify 5n^2-7n?

OpenStudy (anonymous):

thats the simplest form

OpenStudy (anonymous):

That's what I thought.

OpenStudy (anonymous):

n(5n-7)

OpenStudy (anonymous):

Thats factor........ @protosbg not simplest form

OpenStudy (anonymous):

i dont know what is simplest form but i think its n(5n-7) the answer becouse its no sense ....

OpenStudy (anonymous):

What about for this one 12(4+3c). How do you simplify this one?

OpenStudy (anonymous):

=12(4+3c) =48+36c

OpenStudy (anonymous):

now, thats the simplest form for 12(4+3c)